distance 200 m,east time elapsed
300 m,west 400 sec
find for velocity
100 m,east time elapsed
400 m,west 400 sec
1
Expert's answer
2012-11-14T07:13:13-0500
1) 200+300m=500m - full distance
velocity is S/t=500/400=1.25 m/s
2)100+400m=500m
velocity is S/t=1.25 m/s